OpenStudy (anonymous):

Divide. Express your answer in lowest terms. 3 4 _ --- / ( _ --- ) = 14 7

OpenStudy (zpupster):

\[\frac{ -3 }{ 14 }\div \frac{ -4 }{ 7 }\]

OpenStudy (zpupster):

like that??

OpenStudy (anonymous):

yea

OpenStudy (zpupster):

\[\frac{ -3 }{ 14 }\times \frac{ 7 }{ -4 }\]

OpenStudy (anonymous):

?

OpenStudy (zpupster):

you see we inverted the 2nd fraction

OpenStudy (zpupster):

the 7 is now the numerator

OpenStudy (anonymous):

oh

OpenStudy (zpupster):

now just multiply the numerators ( top numbers) and denominators( bottom numbers)

OpenStudy (anonymous):

-21 -56

OpenStudy (zpupster):

a negative divided by a negative is a positive so we can get rid of minus signs also what number goes into 21 and 56

OpenStudy (anonymous):

2?

OpenStudy (zpupster):

21 is odd so 2 can not be divided into 21

OpenStudy (anonymous):

3 or 1? 21 / 3 = 7

OpenStudy (zpupster):

3 does not go into 56 how about 7??

OpenStudy (anonymous):

56 / 7 = 8

OpenStudy (zpupster):

so 7 x 3/7 x 8 the 7's cancel and our answer is n lowest terms. ??

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

u need help :)

OpenStudy (godlovesme):

@AidanWithanAnota

OpenStudy (zpupster):

3/8 remember the 7's cancel

OpenStudy (godlovesme):

nice job :) @zpupster

OpenStudy (zpupster):

ty